# The impact of $\gamma N$ and $\gamma^{*}N$ interactions on our understanding of nucleon excitations

### v1

@@ 6.1.1   Wording
Resonances are unstable particles above thresholds that decay into specific two- or multihadron final states. For example, the $\Delta(1232)\nicefrac{{3}}{{2}}^{+}$ is above the $N\pi$ threshold and decays almost exclusively into $N\pi$. The $\Delta(1232)\nicefrac{{3}}{{2}}^{+}$ wave function thus contains a $N\pi$ component [628] but is dominantly a three-quark state. The Roper resonance $N(1440)\nicefrac{{1}}{{2}}^{+}$ decays into $N\pi$, $N\sigma$, and $\Delta(1232)\nicefrac{{3}}{{2}}^{+}\pi$. Within effective field theory frameworks, the $N(1440)\nicefrac{{1}}{{2}}^{+}$ can be dynamically generated from meson-baryon interactions, and no three-quark component is required. When this mechanism dominates, the resonance is often referred to as a hadronic molecule. A baryon that arises dynamically from meson-baryon interactions must, at the quark level, contain at least four quarks and one antiquark. If such a state is narrow, it is sometimes classified as a pentaquark.
A state is called exotic if either its $J^{PC}$ quantum numbers or its minimal quark content ($q\bar{q}$ for mesons or $qqq$ for baryons) are not compatible with the conventional quark model. Hadrons in which the gluonic field contributes excitation energy are referred to as hybrids. Only mesonic hybrids can have quantum-number-exotic $J^{PC}$ combinations, i.e., values that are forbidden for $q\bar{q}$ configurations in the quark model. Finally, states that are not part of the traditional quark model spectrum but also do not possess exotic quantum numbers are sometimes termed cryptoexotic.
@@
A key challenge is to disentangle the interplay between the intrinsic quark core and the surrounding meson cloud. A useful conceptual tool in this context is Weinberg’s compositeness criterion [966], which relates observable scattering parameters to the internal structure of a bound state. Specifically, Weinberg established a connection between the scattering length $a$, the effective range $r$, and the probability $Z$ of finding a compact, or “elementary", component within the wave function of a hadronic bound-state:
@@
The compositeness criterion has been generalized to account for coupled-channel dynamics [967], though the interpretation of the effective range in such systems remains nontrivial and warrants special attention [968]. In the present context, we do not elaborate on these refinements, but instead highlight recent results obtained within the Jülich-Bonn model [651]. This study tested multiple definitions of compositeness and analyzed 13 light-quark baryon resonances. For eight of these states, reasonably consistent results were found. Four states, the $N(1440)\nicefrac{{1}}{{2}}^{+}$, $N(1520)\nicefrac{{3}}{{2}}^{-}$, $N(1535)\nicefrac{{1}}{{2}}^{-}$, and $N(1710)\nicefrac{{1}}{{2}}^{+}$, “have the chance of being composite". For example, the $N(1440)\nicefrac{{1}}{{2}}^{+}$ appears to have an approximately 50% probability of being “elementary", with the remaining strength arising from molecular components distributed among its dominant decay channels. Obviously, it is the “elementary" part that is responsible for the node in the wave function found in electroproduction, with $N\pi$, $\Delta(1232)\nicefrac{{3}}{{2}}^{+}\pi$, and $N\sigma$ forming the meson cloud. These findings support a nuanced interpretation in which a dynamically generated resonance may still contain a genuine three-quark core, dressed by meson-baryon components. In this view, the $N(1440)\nicefrac{{1}}{{2}}^{+}$ is not dynamically generated in nature but can be dynamically generated within a suitable theoretical framework.
@@
As of today, the picture for the $N(1440)\nicefrac{{1}}{{2}}^{+}$ seems reasonably consistent: it is frequently described as a dynamically generated state, as in Ref. [651], where it includes a three-quark core, and in electroproduction studies, where it emerges as a quark-model state dressed by a meson cloud. In contrast, conflicting interpretations exist for the $N(1520)\nicefrac{{3}}{{2}}^{-}$. In Ref. [651], this resonance is characterized as dominantly ($\sim 75\%$) molecular in nature. However, electroproduction data reveal it as a compact object with minimal meson-cloud contributions.
@@
An important question remains unanswered: what is the underlying reason for the existence of the $N(1440)\nicefrac{{1}}{{2}}^{+}$, or more generally, of dynamically generated states? Are resonances like the $N(1440)\nicefrac{{1}}{{2}}^{+}$ molecular in nature, coupling to a three-quark core via the annihilation of a $q\bar{q}$ pair? Or does a pole, originating from the strong interactions among three quarks, often referred to as a CDD pole in the literature [969], induce the attraction between the decay products?
@@
While this may resemble a chicken-and-egg problem, the implications are substantial. If the strong force is the primary mechanism behind the existence of resonances, then the $\Delta\pi$, $N\rho$, and $N\sigma$ interactions should all resonate at the same mass. Conversely, if molecular interactions are the main driver, a $J^{P}=\nicefrac{{1}}{{2}}^{+}$ state at 1763 MeV coupling to $\Delta\pi$, $N\rho$, and $N\sigma$ could generate three distinct resonances [437]. This latter approach was recently criticized in Ref. [884] for not offering “a meaningful physical understanding of the origin of resonances.” Moreover, it does not fully account for the electroproduction data on the $N(1440)\nicefrac{{1}}{{2}}^{+}$, see Sec. 5.2.2.
@@
$P_{c}(4312)$ $4311.9\pm 0.7^{+6.8}_{-0.6}$ $\Sigma_{c}^{+}\bar{D}^{0}$ 4317.5 $9.8\pm 2.7^{+3.7}_{-4.5}$ [970]
$P_{c}(4337)$ $4337^{+7}_{-4}{}^{+2}_{-2}$ $\Sigma_{c}^{0}\bar{D}^{+}$ 4323.4 $9.8\pm 2.7^{+3.7}_{-4.5}$ [971]
$P_{c}(4380)^{\dagger}$ $4380\pm 8\pm 28$ $\Sigma_{c}^{0}\bar{D}^{+}$ 4323.4 $205\pm 90$ [972]
$P_{c}(4440)$ $4440.3\pm 1.3^{+4.1}_{-4.7}$ $\Sigma_{c}^{+}\bar{D}^{*0}$ 4459.5 $20.6\pm 4.9^{+8.7}_{-10.1}$ [970]
$P_{c}(4457)$ $4457.3\pm 0.6^{+4.1}_{-1.7}$ $\Sigma_{c}^{0}\bar{D}^{*+}$ 4464.0 $6.4\pm 2.0^{+5.7}_{-1.9}$ [970]
$P_{cs}(4338)$ $4338.2\pm 0.7\pm 0.4$ $\Xi_{c}^{0}D^{+}$ 4340.1 $7.0\pm 1.2\pm 1.3$ [973]
$P_{cs}(4459)$ $4458.8\pm 2.9^{+4.7}_{-1.1}$ $\Xi_{c}^{{}^{\prime}+}\bar{D}^{0}$ 4443.0 $17.3\pm 6.5^{+8.0}_{-5.7}$ [974]
$\dagger$ Not confirmed in [970]
Figure 6.1: LHCb data on $\Lambda_{b}\to J/\psi pK^{-}$ [970] with fit from Ref. [975] using a hadronic molecular picture. The vertical dashed lines are thresholds for $\Sigma_{c}\bar{D}$, $\Sigma_{c}^{*}\bar{D}$, and $\Sigma_{c}\bar{D}^{*}$.
@@ 6.1.3   Heavy pentaquarks
The discovery of narrow peaks in the $J/\psi\,p$ invariant mass spectrum observed in $\Lambda_{b}\to J/\psi\,p\,K^{-}$ decays [976, 972] provided clear evidence for states with a minimal quark content beyond that of the conventional quark model. The decay of an excited three-quark nucleon into $J/\psi\,p$ is highly suppressed by the OZI rule; thus, the minimal quark content of the observed narrow peaks must be $c\bar{c}uud$. The most recent results on $J/\psi\,p$ resonances in $\Lambda_{b}\to J/\psi\,p\,K^{-}$ decays were reported in Ref. [970]. An additional state was observed in $\Lambda_{b}\to J/\psi\,p\,\bar{p}$ [971]. Peaks in the $J/\psi\,\Lambda$ invariant mass spectrum have also been seen in the decays $\Xi_{b}^{-}\to J/\psi\,\Lambda\,K^{-}$ [974] and $B^{-}\to J/\psi\,\Lambda\,\bar{p}$ [973], indicating a minimal quark content of $c\bar{c}uds$. Table 6.1 summarizes the current experimental results on heavy pentaquark candidates.
@@
6.1.4   Light-quark pentaquarks
In 1997, Diakonov, Petrov and Polyakov predicted the existence of an anti-decuplet of exotic light-quark pentaquarks based on the chiral soliton model, a theoretical framework in which baryons emerge as topological solitons of the pion field rather than as three-quark states [977]. Translated into quark-model language, this implies that baryons do not consist solely of three quarks, but also include sea quarks. When $q\bar{q}$ pairs are assembled in light chiral fields, their production incurs little additional energy, provided that their flavor content matches the quantum numbers of the baryon. The lightest state in the anti-decuplet, denoted $\Theta^{+}$, was predicted to have strangeness $S=+1$, spin-parity $J=\nicefrac{{1}}{{2}}^{+}$, a mass of 1530 MeV, and to be narrow in width. Its minimal quark content is given by $\Theta^{+}=uud\,d\bar{s}$, identifying it as an exotic pentaquark. The $\Theta^{+}$ should be accompanied by a nucleon-like isospin doublet, three $\Sigma$-like states, and a quartet of $\Xi$ states, among which $\Xi^{--}$ and $\Xi^{+}$ are exotic.
The 2003 announcement of the discovery of the $\Theta^{+}(1540)$ in the reaction $\gamma\,n\to K^{+}K^{-}n$ on a carbon target [247] generated tremendous excitement in both the medium-energy and high-energy physics communities. Within less than a year, the initial observation was supported by similar results from ten other experiments. Shortly thereafter, a peak in the $\Xi^{-}\pi^{-}$ invariant mass spectrum was reported by the NA49 collaboration at CERN and interpreted as the $\Xi^{--}(1862)$ pentaquark [978]. One year later, data from the GRAAL experiment revealed a structure in the $N\eta$ invariant mass distribution, which was interpreted as a $N(1680)$ pentaquark [979]. In the context of these pentaquark searches, two additional candidates were proposed that were not predicted by the chiral soliton model: a doubly charged isovector companion of the $\Theta^{+}$, and a narrow anti-charmed baryon denoted $\Theta_{c}$.
[…]
Detailed discussions of the experiments that claimed evidence for pentaquark candidates, as well as those that reported null results, can be found in a number of reviews [980, 981, 982, 212, 983, 984, 985, 986, 987, 988, 989]. The Particle Data Group ultimately concluded that the non-existence of pentaquarks appears compelling [990, 991]. While the chiral soliton model offered an elegant and appealing picture of the baryon structure, it did not withstand experimental scrutiny.
@@ 6.2   The $\Lambda(1405)$ and its debated two-pole structure
The $\Lambda(1405)$ resonance was first observed in a bubble chamber experiment at the Lawrence Radiation Laboratory in 1961 [992]. Tripp et al. [993] determined the relative signs of $\bar{K}N\to\pi\Sigma$ transition amplitudes for the $\Sigma(1385)$, $\Lambda(1405)$, and $\Lambda(1520)$, identifying the $\Lambda(1405)$ as primarily an $SU(3)_{f}$ singlet state. In Ref. [994], the $\Lambda(1405)$ was investigated in the production and decay chain $K^{-}p\to\pi^{-}\Sigma^{+}(1670)\nicefrac{{3}}{{2}}^{-}$, $\Sigma^{+}(1670)\nicefrac{{3}}{{2}}^{-}\to\Lambda(1405)\pi^{+}$, $\Lambda(1405)\to\Sigma^{\pm}\pi^{\mp}$. The sign of the transition amplitude enables a determination of the $\Lambda(1405)$ $SU(3)_{f}$ structure [995]. Later experimental studies mainly focused on measuring the line shape of $\Lambda(1405)$ [996, 997, 998, 999, 1000, 186, 1001, 1002, 1003, 1004, 1005, 239, 1006] or confirming its spin and parity [178]. Recent investigations of the reactions $pp\to\Sigma^{-}\pi^{+}pK^{+}$[1003] and $\gamma p\to K^{+}\Sigma^{0}\pi^{0}$ [239, 1006] provided evidence for a two-pole structure of the invariant mass distribution $\Sigma\pi$ in the $\Lambda(1405)$ region.
@@
Concerning its theoretical interpretations, Dalitz and collaborators already predicted the existence of the $\Lambda(1405)$ as a quasibound molecular state of the $\bar{K}N$ system [1007, 1008] two years before its discovery based on the $\bar{K}N$ scattering lengths [1009]. Later, Kaiser, Waas and Weise constructed an effective potential from a chiral Lagrangian, and the $\Lambda(1405)$ emerged as a quasi-bound state in the $\bar{K}N$ and $\pi\Sigma$ coupled-channel system [1010, 1011]. In quark models, the $\Lambda(1405)$ and $\Lambda(1520)$ are interpreted as $qqq$ resonances where one of the quarks is excited to the $p$ state, forming a spin-doublet of states with a dominant $SU(3)_{f}$ singlet structure [535]. This is also the interpretation in later quark-model calculations of baryon resonances [548, 666, 677]. However, such interpretations need not necessarily conflict; resonances may have a dual nature. For example, the Roper resonance $N(1440)\nicefrac{{1}}{{2}}^{+}$ has an unexpectedly low mass of about 100 MeV below the $N(1535)\nicefrac{{1}}{{2}}^{-}$, while quark models predict the mass of the nucleon’s first radial excitation above its parity partner. This discrepancy led to non-$qqq$ interpretations of the Roper resonance, and a large number of papers identified it as a molecular state from very early times [1012] to recent publications [651] (although the $N(1440)\nicefrac{{1}}{{2}}^{+}$ has, in the latter publication, a significant $qqq$ contribution). In electroproduction, the Roper resonance reveals a $qqq$ core surrounded by a meson cloud [546]. Hence, interpretations of states as dynamically generated resonances and as quark-model states can co-exist without necessarily being in conflict.
@@
Oller and Meißner studied the S-wave kaon–nucleon interactions for strangeness $S=-1$ in a relativistic chiral unitary approach based on coupled channels [1013]. The construction of meson-baryon potentials compatible with chiral symmetry generated two poles in the $\Lambda(1405)$ region. This exciting finding initiated a large number of further investigations [1014, 624, 1015, 1016, 1017, 1018, 1019, 1020, 1021, 1022, 1023, 1024, 1025, 1026, 1027, 1028, 1029, 1030, 1031, 605], and several reviews have highlighted the nature of the $\Lambda(1405)$ [1032, 606, 1033, 602, 603, 1034]. The resulting three poles $\Lambda(1380)$, $\Lambda(1405)$, and $\Lambda(1680)$ are suggested to be the singlet and the two octet isoscalar states, as expected from the decomposition $8\otimes 8=1\oplus 8\oplus 8\oplus 10\oplus\overline{10}\oplus 27$. Figure 6.2 shows how low-mass strange baryons with $J^{P}=\nicefrac{{1}}{{2}}^{-}$ emerge when the $SU(3)_{f}$ breaking coefficient is tuned from 0 (the $SU(3)_{f}$ limit) to 1 (the physical point).

### v4

@@ 6.1.1  Wording
Resonances are unstable particles above thresholds that decay into specific two- or multihadron final states. For example, the $\Delta(1232)\nicefrac{{3}}{{2}}^{+}$ is above the $N\pi$ threshold and decays almost exclusively into $N\pi$. The $\Delta(1232)\nicefrac{{3}}{{2}}^{+}$ wave function thus contains a $N\pi$ component Pascalutsa:2006up but is dominantly a three-quark state. The Roper resonance $N(1440)\nicefrac{{1}}{{2}}^{+}$ decays into $N\pi$, $N\sigma$, and $\Delta(1232)\nicefrac{{3}}{{2}}^{+}\pi$. Within effective field theory frameworks, the $N(1440)\nicefrac{{1}}{{2}}^{+}$ can be dynamically generated from meson-baryon interactions, and no three-quark component is required. When this mechanism dominates, the resonance is often referred to as a hadronic molecule. A baryon that arises dynamically from meson-baryon interactions must, at the quark level, contain at least four quarks and one antiquark. If such a state is narrow, it is sometimes classified as a pentaquark.
A state is called exotic if either its $J^{PC}$ quantum numbers or its minimal quark content ($q\bar{q}$ for mesons or qqqqqq for baryons) are not compatible with the conventional quark model. Hadrons in which the gluonic field contributes excitation energy are referred to as hybrids. Only mesonic hybrids can have quantum-number-exotic $J^{PC}$ combinations, i.e., values that are forbidden for $q\bar{q}$ configurations in the quark model. Finally, states that are not part of the traditional quark model spectrum but also do not possess exotic quantum numbers are sometimes termed cryptoexotic.
@@
A key challenge is to disentangle the interplay between the intrinsic quark core and the surrounding meson cloud. A useful conceptual tool in this context is Weinberg’s compositeness criterion Weinberg:1965zz , which relates observable scattering parameters to the internal structure of a bound state. Specifically, Weinberg established a connection between the scattering length $a$, the effective range $r$, and the probability $Z$ of finding a compact, or “elementary", component within the wave function of a hadronic bound-state:
@@
The compositeness criterion has been generalized to account for coupled-channel dynamics Matuschek:2020gqe , though the interpretation of the effective range in such systems remains nontrivial and warrants special attention Baru:2021ldu . In the present context, we do not elaborate on these refinements, but instead highlight recent results obtained within the Jülich-Bonn model Wang:2023snv . This study tested multiple definitions of compositeness and analyzed 13 light-quark baryon resonances. For eight of these states, reasonably consistent results were found. Four states, the $N(1440)\nicefrac{{1}}{{2}}^{+}$, $N(1520)\nicefrac{{3}}{{2}}^{-}$, $N(1535)\nicefrac{{1}}{{2}}^{-}$, and $N(1710)\nicefrac{{1}}{{2}}^{+}$, “have the chance of being composite". For example, the $N(1440)\nicefrac{{1}}{{2}}^{+}$ appears to have an approximately 50% probability of being “elementary", with the remaining strength arising from molecular components distributed among its dominant decay channels. Obviously, it is the “elementary" part that is responsible for the node in the wave function found in electroproduction, with $N\pi$, $\Delta(1232)\nicefrac{{3}}{{2}}^{+}\pi$, and $N\sigma$ forming the meson cloud. These findings support a nuanced interpretation in which a dynamically generated resonance may still contain a genuine three-quark core, dressed by meson-baryon components. In this view, the $N(1440)\nicefrac{{1}}{{2}}^{+}$ is not dynamically generated in nature but can be dynamically generated within a suitable theoretical framework.
@@
As of today, the picture for the $N(1440)\nicefrac{{1}}{{2}}^{+}$ seems reasonably consistent: it is frequently described as a dynamically generated state, as in Ref. Wang:2023snv , where it includes a three-quark core, and in electroproduction studies, where it emerges as a quark-model state dressed by a meson cloud. In contrast, conflicting interpretations exist for the $N(1520)\nicefrac{{3}}{{2}}^{-}$. In Ref. Wang:2023snv , this resonance is characterized as dominantly ($\sim 75\%$) molecular in nature. However, electroproduction data reveal it as a compact object with minimal meson-cloud contributions.
@@
An important question remains unanswered: what is the underlying reason for the existence of the $N(1440)\nicefrac{{1}}{{2}}^{+}$, or more generally, of dynamically generated states? Are resonances like the $N(1440)\nicefrac{{1}}{{2}}^{+}$ molecular in nature, coupling to a three-quark core via the annihilation of a $q\bar{q}$ pair? Or does a pole, originating from the strong interactions among three quarks, often referred to as a CDD pole in the literature Castillejo:1955ed , induce the attraction between the decay products?
@@
While this may resemble a chicken-and-egg problem, the implications are substantial. If the strong force is the primary mechanism behind the existence of resonances, then the $\Delta\pi$, $N\rho$, and $N\sigma$ interactions should all resonate at the same mass. Conversely, if molecular interactions are the main driver, a $J^{P}=\nicefrac{{1}}{{2}}^{+}$ state at 1763 MeV coupling to $\Delta\pi$, $N\rho$, and $N\sigma$ could generate three distinct resonances Suzuki:2009nj . This latter approach was recently criticized in Ref. Owa:2025mep for not offering “a meaningful physical understanding of the origin of resonances.” Moreover, it does not fully account for the electroproduction data on the $N(1440)\nicefrac{{1}}{{2}}^{+}$, see Sec. 5.2.2.
@@
$P_{c}(4312)$ $4311.9\pm 0.7^{+6.8}_{-0.6}$ $\Sigma_{c}^{+}\bar{D}^{0}$ 4317.5 $9.8\pm 2.7^{+3.7}_{-4.5}$ LHCb:2019kea
$P_{c}(4337)$ $4337^{+7}_{-4}{}^{+2}_{-2}$ $\Sigma_{c}^{0}\bar{D}^{+}$ 4323.4 $9.8\pm 2.7^{+3.7}_{-4.5}$ LHCb:2021chn
$P_{c}(4380)^{\dagger}$ $4380\pm 8\pm 28$ $\Sigma_{c}^{0}\bar{D}^{+}$ 4323.4 $205\pm 90$ LHCb:2016ztz
$P_{c}(4440)$ $4440.3\pm 1.3^{+4.1}_{-4.7}$ $\Sigma_{c}^{+}\bar{D}^{*0}$ 4459.5 $20.6\pm 4.9^{+8.7}_{-10.1}$ LHCb:2019kea
$P_{c}(4457)$ $4457.3\pm 0.6^{+4.1}_{-1.7}$ $\Sigma_{c}^{0}\bar{D}^{*+}$ 4464.0 $6.4\pm 2.0^{+5.7}_{-1.9}$ LHCb:2019kea
$P_{cs}(4338)$ $4338.2\pm 0.7\pm 0.4$ $\Xi_{c}^{0}D^{+}$ 4340.1 $7.0\pm 1.2\pm 1.3$ LHCb:2022ogu
$P_{cs}(4459)$ $4458.8\pm 2.9^{+4.7}_{-1.1}$ $\Xi_{c}^{{}^{\prime}+}\bar{D}^{0}$ 4443.0 $17.3\pm 6.5^{+8.0}_{-5.7}$ LHCb:2020jpq
$\dagger$ Not confirmed in LHCb:2019kea
Figure 6.1: LHCb data on $\Lambda_{b}\to J/\psi pK^{-}$ LHCb:2019kea with fit from Ref. Shen:2024nck using a hadronic molecular picture. The vertical dashed lines are thresholds for $\Sigma_{c}\bar{D}$, $\Sigma_{c}^{*}\bar{D}$, and $\Sigma_{c}\bar{D}^{*}$.
@@ 6.1.3  Heavy pentaquarks
The discovery of narrow peaks in the $J/\psi\,p$ invariant mass spectrum observed in $\Lambda_{b}\to J/\psi\,p\,K^{-}$ decays LHCb:2015yax , LHCb:2016ztz provided clear evidence for states with a minimal quark content beyond that of the conventional quark model. The decay of an excited three-quark nucleon into $J/\psi\,p$ is highly suppressed by the OZI rule; thus, the minimal quark content of the observed narrow peaks must be $c\bar{c}uud$. The most recent results on $J/\psi\,p$ resonances in $\Lambda_{b}\to J/\psi\,p\,K^{-}$ decays were reported in Ref. LHCb:2019kea . An additional state was observed in $\Lambda_{b}\to J/\psi\,p\,\bar{p}$ LHCb:2021chn . Peaks in the $J/\psi\,\Lambda$ invariant mass spectrum have also been seen in the decays $\Xi_{b}^{-}\to J/\psi\,\Lambda\,K^{-}$ LHCb:2020jpq and $B^{-}\to J/\psi\,\Lambda\,\bar{p}$ LHCb:2022ogu , indicating a minimal quark content of $c\bar{c}uds$. Table 6.1 summarizes the current experimental results on heavy pentaquark candidates.
There are numerous interpretations of these structures. Frequently, the observed peaks are described as loosely bound molecular states composed of charmed baryons and anti-charmed mesons. In Fig. 6.1, we show the LHCb data on $\Lambda_{b}\to J/\psi\,p\,K^{-}$ LHCb:2019kea , along with a fit from Ref. Shen:2024nck based on a hadronic molecular model. The small structure near 4380 MeV, originally proposed as a broad resonance in Refs. LHCb:2015yax , LHCb:2016ztz but not confirmed in later LHCb data LHCb:2019kea , is now described by a pole at (4366.5i 4.7)(4366.5-i\,4.7) MeV. Two alternative fits yield slightly different values. Some of these states may originate from charmonium bound to light-quark baryons. Alternatively, they could arise from threshold singularities due to final-state rescattering, which leads to a logarithmic branch point in the amplitude. Other interpretations suggest compact states formed from triquark–diquark configurations, or systems in which a $c\bar{c}$ core is surrounded by light quarks. QCD sum rules have also been employed to describe their properties. Several comprehensive reviews on tetraquarks, pentaquarks, and their possible interpretations are available; see Refs. Brambilla:2019esw , Liu:2019zoy , Dong:2021bvy , Dong:2021juy , Chen:2022asf , Burns:2022uiv , Klempt:2022zwo , Liu:2024uxn .
6.1.4  Light-quark pentaquarks
In 1997, Diakonov, Petrov and Polyakov predicted the existence of an anti-decuplet of exotic light-quark pentaquarks based on the chiral soliton model, a theoretical framework in which baryons emerge as topological solitons of the pion field rather than as three-quark states Diakonov:1997mm . Translated into quark-model language, this implies that baryons do not consist solely of three quarks, but also include sea quarks. When $q\bar{q}$ pairs are assembled in light chiral fields, their production incurs little additional energy, provided that their flavor content matches the quantum numbers of the baryon. The lightest state in the anti-decuplet, denoted $\Theta^{+}$, was predicted to have strangeness $S=+1$, spin-parity $J=\nicefrac{{1}}{{2}}^{+}$, a mass of 1530 MeV, and to be narrow in width. Its minimal quark content is given by $\Theta^{+}=uud\,d\bar{s}$, identifying it as an exotic pentaquark. The $\Theta^{+}$ should be accompanied by a nucleon-like isospin doublet, three $\Sigma$-like states, and a quartet of $\Xi$ states, among which $\Xi^{--}$ and $\Xi^{+}$ are exotic.
The 2003 announcement of the discovery of the $\Theta^{+}(1540)$ in the reaction $\gamma\,n\to K^{+}K^{-}n$ on a carbon target LEPS:2003wug generated tremendous excitement in both the medium-energy and high-energy physics communities. Within less than a year, the initial observation was supported by similar results from ten other experiments. Shortly thereafter, a peak in the $\Xi^{-}\pi^{-}$ invariant mass spectrum was reported by the NA49 collaboration at CERN and interpreted as the $\Xi^{--}(1862)$ pentaquark NA49:2003fxh . One year later, data from the GRAAL experiment revealed a structure in the $N\eta$ invariant mass distribution, which was interpreted as a $N(1680)$ pentaquark GRAAL:2004ndn . In the context of these pentaquark searches, two additional candidates were proposed that were not predicted by the chiral soliton model: a doubly charged isovector companion of the $\Theta^{+}$, and a narrow anti-charmed baryon denoted $\Theta_{c}$.
@@ 6.1.4  Light-quark pentaquarks
Detailed discussions of the experiments that claimed evidence for pentaquark candidates, as well as those that reported null results, can be found in a number of reviews Klempt:2004yz , Kabana:2005tp , Danilov:2005kt , Burkert:2005ft , Danilov:2008uxa , Hicks:2012zz , Liu:2014yva , Strakovsky:2024ppo , Kim:2024tae , Praszalowicz:2024mji , Praszalowicz:2024zsy . The Particle Data Group ultimately concluded that the non-existence of pentaquarks appears compelling ParticleDataGroup:2006fqo , ParticleDataGroup:2008zun . While the chiral soliton model offered an elegant and appealing picture of the baryon structure, it did not withstand experimental scrutiny.
@@ 6.2  The $\Lambda(1405)$ and its debated two-pole structure
The $\Lambda(1405)$ resonance was first observed in a bubble chamber experiment at the Lawrence Radiation Laboratory in 1961 Alston:1961zzd . Tripp et al. Tripp:1968ukt determined the relative signs of $\bar{K}N\to\pi\Sigma$ transition amplitudes for the $\Sigma(1385)$, $\Lambda(1405)$, and $\Lambda(1520)$, identifying the $\Lambda(1405)$ as primarily an $SU(3)_{f}$ singlet state. In Ref. Hemingway:1984pz , the $\Lambda(1405)$ was investigated in the production and decay chain $K^{-}p\to\pi^{-}\Sigma^{+}(1670)\nicefrac{{3}}{{2}}^{-}$, $\Sigma^{+}(1670)\nicefrac{{3}}{{2}}^{-}\to\Lambda(1405)\pi^{+}$, $\Lambda(1405)\to\Sigma^{\pm}\pi^{\mp}$. The sign of the transition amplitude enables a determination of the $\Lambda(1405)$ $SU(3)_{f}$ structure Klempt:2020klf . Later experimental studies mainly focused on measuring the line shape of $\Lambda(1405)$ Zychor:2007mm , Esmaili:2009rf , Moriya:2009mx , Moriya:2010zz , HADES:2012csk , CLAS:2013rjt , CLAS:2013zie , Ahn:2010zzb , Siebenson:2013rpa , Ren:2015bsa , J-PARCE31:2022plu , BGOOD:2021sog , Wickramaarachchi:2022mhi or confirming its spin and parity CLAS:2014tbc . Recent investigations of the reactions $pp\to\Sigma^{-}\pi^{+}pK^{+}$Siebenson:2013rpa and $\gamma p\to K^{+}\Sigma^{0}\pi^{0}$ BGOOD:2021sog , Wickramaarachchi:2022mhi provided evidence for a two-pole structure of the invariant mass distribution $\Sigma\pi$ in the $\Lambda(1405)$ region.
@@
Concerning its theoretical interpretations, Dalitz and collaborators already predicted the existence of the $\Lambda(1405)$ as a quasibound molecular state of the $\bar{K}N$ system Dalitz:1960du , Dalitz:1967fp two years before its discovery based on the $\bar{K}N$ scattering lengths Dalitz:1959dn . Later, Kaiser, Waas and Weise constructed an effective potential from a chiral Lagrangian, and the $\Lambda(1405)$ emerged as a quasi-bound state in the $\bar{K}N$ and $\pi\Sigma$ coupled-channel system Kaiser:1995eg , Kaiser:1996js . In quark models, the $\Lambda(1405)$ and $\Lambda(1520)$ are interpreted as qqqqqq resonances where one of the quarks is excited to the $p$ state, forming a spin-doublet of states with a dominant $SU(3)_{f}$ singlet structure Isgur:1978xj . This is also the interpretation in later quark-model calculations of baryon resonances Capstick:1986ter , Glozman:1997ag , Loring:2001ky . However, such interpretations need not necessarily conflict; resonances may have a dual nature. For example, the Roper resonance $N(1440)\nicefrac{{1}}{{2}}^{+}$ has an unexpectedly low mass of about 100 MeV below the $N(1535)\nicefrac{{1}}{{2}}^{-}$, while quark models predict the mass of the nucleon’s first radial excitation above its parity partner. This discrepancy led to non-qqqqqq interpretations of the Roper resonance, and a large number of papers identified it as a molecular state from very early times Bhargava:1969jt to recent publications Wang:2023snv (although the $N(1440)\nicefrac{{1}}{{2}}^{+}$ has, in the latter publication, a significant qqqqqq contribution). In electroproduction, the Roper resonance reveals a qqqqqq core surrounded by a meson cloud Burkert:2017djo . Hence, interpretations of states as dynamically generated resonances and as quark-model states can co-exist without necessarily being in conflict.
@@
Oller and Meißner studied the S-wave kaon–nucleon interactions for strangeness S=1S=-1 in a relativistic chiral unitary approach based on coupled channels Oller:2000fj . The construction of meson-baryon potentials compatible with chiral symmetry generated two poles in the $\Lambda(1405)$ region. This exciting finding initiated a large number of further investigations Oset:2001cn , Lutz:2001yb , Jido:2002zk , Garcia-Recio:2002yxy , Jido:2003cb , Oller:2006jw , Doring:2010rd , Haidenbauer:2010ch , Mai:2014xna , Feijoo:2015yja , Cieply:2015pwa , Cieply:2016jby , Liu:2016wxq , Ramos:2016odk , Kamiya:2016jqc , Miyahara:2018onh , Bruns:2021krp , Lu:2022hwm , Sadasivan:2022srs , Guo:2023wes , and several reviews have highlighted the nature of the $\Lambda(1405)$ Hyodo:2011ur , Guo:2017jvc , Hyodo:2020czb , Meissner:2020khl , Mai:2020ltx , Nieves:2024dcz . The resulting three poles $\Lambda(1380)$, $\Lambda(1405)$, and $\Lambda(1680)$ are suggested to be the singlet and the two octet isoscalar states, as expected from the decomposition $8\otimes 8=1\oplus 8\oplus 8\oplus 10\oplus\overline{10}\oplus 27$. Figure 6.2 shows how low-mass strange baryons with $J^{P}=\nicefrac{{1}}{{2}}^{-}$ emerge when the $SU(3)_{f}$ breaking coefficient is tuned from 0 (the $SU(3)_{f}$ limit) to 1 (the physical point), see also Ref. Zhuang:2024udv .